solve for b where b^4 +13b^2 + 36 = 0
Answer by mananth(16946)

b^4 +13b^2 + 36 = 0
b^4+9b^2+4b^2+36=0
b^2(b^2+9)+4(b^2+9)=0
(b^2+9)(b^2+4)=0
...
m.ananth@hotmail.ca